# Convergent Privacy Framework for Multi-layer GNNs through Contractive Message Passing

### v2

@@ III-A2 Empirical (Counter-Intuitive) Observation: Deeper GNNs May Enhance Privacy
Recent work by [wu2024provable] revealed a counter-intuitive phenomenon: deeper GNNs empirically exhibit lower vulnerability to membership and link inference attacks. This challenges standard privacy composition analysis, which suggests privacy risk increases with model depth due to multiple queries of the graph data. The insight stems from the over-smoothing [^1] Incidentally, another similar phrase is “over-fitting”, which refers to a model performing well on training data but poorly on unseen data due to memorization. We emphasize that they are different concepts to avoid confusion. [aaai/ChenLLLZS20] phenomenon, where node representations become indistinguishably homogeneous phenomenon, where node representations converge toward similar values as depth increases, making it inherently difficult for adversaries to distinguish individual nodes or infer sensitive relationships.
@@ III-B Core Idea for Convergent Privacy
Designing Private GNNs with Convergent Privacy. We aim to leverage the privacy amplification properties of contractive map to design a new framework for private GNNs. This framework is inspired by the recent advances in differentially private gradient descent (DP-GD) [nips/ChourasiaYS21, icml/BokShifted24], which has shown that the privacy cost can converge to a finite value even with arbitrarily many iterations. Motivated by it, we aim to translate the advanced privacy analysis techniques from DP-GD to GNNs. We observe that perturbed message passing (Equation 2) in GNNs follows a strikingly mathematical parallel pattern as DP-GD: This parallel structure enables to leverage the insight of convergent privacy analysis in the context of GNNs, provided we ensure two critical conditions:
@@ IV Private GNNs with Contractiveness
Security Model. Aggregation-based GNNs such as GCN [iclr/kipf2017semi], GCN-II [icml/ChenWHDL20], and SAGE [nips/HamiltonYL17] reveal only the final node embeddings, keeping intermediate embeddings hidden. The final node embeddings learned in this DP manner can be subsequently applied to downstream tasks such as node classification [www/WangLDLW24], link prediction [www/RaoKR24], and graph classification [icdm/LiW024]. Since DP introduces utility loss, GAP [uss/sajadmanesh2023gap] seeks to improve model utility by concatenating all intermediate node embeddings from the $K$ layers. However, it causes the noise scale $\sigma$ to grow linearly as $O(\sqrt{{K}/{\epsilon^{2}}})$, and this privacy bound is loose. This work takes a step further beyond GAP, removing the assumption that all intermediate node embeddings must be revealed under a more realistic security model [nips/0001S22].
@@ IV-A Contractive Aggregation Module
Contractive operations in GNNs, such as the graph convolution (GConv) layer [iclr/KipfW17], inherently reduce privacy risks by mitigating the memorization of GNN parameters through the over-smoothing phenomenon [aaai/ChenLLLZS20]. This property aligns well with the need for $K$-layer aggregation in long-range graphs, where a target node aggregates embeddings from distant source nodes. However, directly stacking $K$ GConv layers, as shown in Figure 3, is suboptimal due to limited expressive power [icml/ChenWHDL20] and heightened sensitivity to DP noise [WuSZFYW19].
@@
Comparison with existing design. GAP [uss/sajadmanesh2023gap] enhances expressiveness by connecting all intermediate embeddings but lacks contractiveness, compromising utility and privacy. In contrast, CGL introduces adjustable coefficients, achieving a balanced trade-off between privacy, utility, and generalizability. Table I summarizes the general design characteristics between GAP and $\mathsf{CARIBOU}$.
@@ IV-C Privacy Auditing Module
Message passing mechanisms integrate graph structures by recursively aggregating and transforming node features from neighbors. Membership Inference Attacks (MIA) on graph examine the vulnerability of message passing to infer whether specific nodes or edges were part of a GNN’s training set [tpsisa/OlatunjiNK21]. The adversaries exploit black-box access to the GNN, querying it with selected data and analyzing outputs (e.g., class probabilities) to infer membership. To align theoretical privacy guarantees with practical deployments, we propose an empirical auditing module. This module simulates an adversary to evaluate GNN privacy before deployment. Extending Carlini et al. [sp/CarliniCN0TT22]’s MIA framework to graphs, we define a graph-specific privacy auditing game (Definition 7) and implement two real-world attacks [sp/Wulinkteller22, tpsisa/OlatunjiNK21] for privacy verification.
@@ IV-D Putting it Together
To improve classification accuracy, $\mathsf{CARIBOU}$ adopts a head MLP architecture proposed by Sajadmanesh et al. [uss/sajadmanesh2023gap] as the CM. This design ensures that the CM effectively combines the information from both feature sets, enabling robust node classification. Furthermore, the CM guarantees efficient training by leveraging the graph-agnostic features ${\bm{X}}^{(0)}$, ensuring a lower-bound performance even in scenarios where graph topology is unavailable.
@@ V Convergent Privacy Analysis
In this section, we present a convergent privacy analysis for perturbed message-passing GNNs with respect to the number of hops. In Appendix A-A, we review the standard privacy analysis for a one-layer perturbed message-passing GNN, and then illustrate that the privacy cost grows linearly with the number of layers under standard composition theorems. We then shift our focus to a convergent privacy analysis for perturbed GNNs whose message-passing layers are contractive. In particular, we draw upon the framework of contractive noisy iteration (CNI) from [icml/BokShifted24], recasting the multi-layer perturbed GNN as a CNI process. Our analysis reveals that, under hidden intermediate states and contractive message-passing layers, the privacy cost converges as the number of hops increases. Finally, we specialize this result to our proposed $\mathsf{CARIBOU}$. We show that $\mathsf{CARIBOU}$’s message-passing operation is contractive, derive its sensitivity, and thereby establish concrete bounds on both edge-level and node-level differential privacy for arbitrarily many hops. All proofs are deferred to the Appendices A-B,A-C.
@@ V-A Contractive Noisy Iteration and Convergent Privacy
Many GNNs, such as GCN, use "mean-type" aggregation, mixing a node’s representation with that of its neighbors. Intuitively, iterative mixing could "amplify" privacy, but existing analyses yield only linear compositions. Our key observation is that perturbed message passing in a contractive GNN layer behaves analogously to noisy gradient descent or noisy iterative maps [nips/ChourasiaYS21, nips/AltschulerT22, siamcomp/AltschulerBT24], where recent work has demonstrated privacy amplification via iteration.
Below, we introduce the framework of contractive noisy iteration (CNI) from [icml/BokShifted24], and the meta theorem proved by [icml/BokShifted24] that provides a tight privacy guarantee for CNI processes.
Definition 8 (Contractive Noisy Iteration (CNI)[icml/BokShifted24, Definition 3.1]).
@@
A special case of CNI considered in [icml/BokShifted24] is the noisy gradient descent, where the contractive functions are the gradient update steps for some fixed loss function $f$, and the noise distribution $\xi_{k}$ is Gaussian. This situation is similar to the perturbed message passing mechanism in a contractive GNN layer, where the contractive function $\phi_{k}\equiv\mathrm{MP}_{G}$ is the message passing operation for a fixed graph $\bm{G}$}.
The complete privacy analysis of CNI processes in [icml/BokShifted24] involves concepts of trade-off functions (Definition 9) and Gaussian differential privacy (GDP, Definition 10), which we briefly review in Appendix A. GDP can be converted to the more familiar Rényi Differential Privacy (RDP) framework:
Lemma 1 (GDP Implies RDP [icml/BokShifted24, Lemma A.4]).
@@
With this connection established, we now state the key meta theorem from [icml/BokShifted24] that analyzes the CNI process and provides a tight privacy guarantee.
Theorem 2 (Meta Theorem on CNI [icml/BokShifted24, Theorem C.5]).
@@
The theorem above slightly generalizes the original result from [icml/BokShifted24] by relaxing the Lipschitz condition to require $\gamma<1$ only for $k\geq 2$ rather than for all iterations. This relaxation is critical for analyzing our $\mathsf{CARIBOU}$ architecture, where the first message-passing layer includes the residual term $\beta{\bm{X}}^{(0)}$, potentially making it non-contractive while subsequent layers remain contractive. The proof extends the original argument by carefully tracking the influence of the first layer on the privacy guarantee. For completeness, we provide the full proof in Appendices A-B,A-C.
@@
Theorem 3 ([icml/BokShifted24, Theorem 4.2] adapted for GNNs).
@@ VI Experimental Evaluation
Datasets. $\mathsf{CARIBOU}$ was tested over nine graph datasets. Five of the datasets have been broadly used to evaluate GNN methods, including Photo and Computers [corr/abs-1811-05868], Cora and PubMed [sen2008collective], Facebook [corr/abs-1102-2166]. We also adjusted the synthetic chain-structured dataset developed under [nips/GuC0SG20] into various scales, termed Chain-S, Chain-M, Chain-L and Chain-X. The chain-structured dataset has been used to understand the relations between privacy/utility and hops, as described in Section III-A. It is considered as an important benchmark to assist the development of long-range interaction GNNs by the ML community [nips/DwivediRGPWLB22]. More details on datasets, model configurations, and privacy configurations are specialized in Appendix A-D1.
Baselines. We compare $\mathsf{CARIBOU}$ with multiple baselines [nips/chien2023differentially, uss/sajadmanesh2023gap, sp/0011L0022, arxiv/DaigavaneNode21, ccs/KolluriBHS22] about edge-level private GNNs and vanilla GNNs. To our best knowledge, GAP [uss/sajadmanesh2023gap] and DPDGC [nips/chien2023differentially] are the strongest baselines for perturbed message passing under Gaussian mechanisms in both edge/node-level DP GNNs. In addition, we consider another research direction, i.e., first perturbing the graph through randomized response and then training GNNs over the perturbed graph PertGraph [sp/0011L0022]. For a comprehensive evaluation, we adopt both research lines of works as our baselines. MLP is a baseline commonly compared with GNNs to demonstrate how well GNNs utilize graph structures, as it provides a reference counterpart to which GNNs learn the only node features without graph topology. More details of the baseline methods are provided in Appendix A-D3.
@@ VI-B Privacy Auditing
Following the PAM outlined in Section IV-C, we focus on black-box, membership-based privacy audits that match the theoretical guarantees of our DP mechanisms and those of prior perturbed message-passing methods [nips/chien2023differentially, uss/sajadmanesh2023gap]. Under this threat model, LinkTeller [sp/Wulinkteller22] and G-MIA [tpsisa/OlatunjiNK21] are canonical and complementary: LinkTeller targets edge-level membership by asking whether a specific edge exists in the training graph; G-MIA targets node-level membership by deciding whether a node and its connected edges were used during training. We adopt G-MIA’s attacking settings of TSTF, where models have been trained on subgraph and tested on full graph. The adversary knows the whole graph $\bm{G}$} and all edges contained in $\bm{G}$} but has no access to the subgraphs used for early training. Both attacks (i) are specifically designed for GNNs, (ii) operate in the transductive setting considered in our analysis, (iii) require only query access to GNN models, and (iv) are publicly available and already used to evaluate DP-GNN defenses. This makes them ideal choices as mechanism-level auditing tools in $\mathsf{CARIBOU}$.
In Table XI (in Appendix), we report the AUC (Area Under the Curve) score about the attack effectiveness, when $\mathsf{CARIBOU}$ is being attacked. AUC is a major metric to evaluate the membership inference attack [sp/Wulinkteller22, he2021stealing]. Specifically, AUC measures true positive rate against the false positive rate on various classification thresholds, and a score of $0.5$ suggesting random guessing. We found $\mathsf{CARIBOU}$ is very effective against LinkTeller, by dropping the attack AUC from between 0.86 to 0.998 across all standard datasets ($\epsilon=\inf$) to less than 0.500 ($\epsilon$ ranges from 1 to 32). The similar effect was also observed by Wu et al. (e.g., less than 0.5 attack precision for 3-layer GCN for high density belief, shown in Table IX) [sp/Wulinkteller22] and Tang et al. (e.g., less than 0.5 attack AUC sometimes in Figure 10) [popets/TangNAA24]. For G-MIA, its AUC on the 5 datasets are already lower than LinkTeller by a notable margin when $\epsilon=\inf$ (between 0.567 to 0.702 for the 5 datasets), so the impact of $\mathsf{CARIBOU}$ is relatively small. But we observe on Cora, $\mathsf{CARIBOU}$ is able to drop AUC from 0.645 to 0.500.
@@ VIII Related Works
Multi-layer GNNs. Recent literature shows that multi-layer GNNs hold significant potential for modeling long-range dependencies and complex relational structures crucial for many real-world applications. Node labels and attributes may depend on distant nodes, necessitating the aggregation of information over larger receptive fields [tnn/WuPCLZY21] through multi-layer GNNs. Notably, Li et al. [icml/Li0GK21] demonstrated, through the 1000-layer GNN, that increasing the network depth attains substantial gains in accuracy, e.g., from $72\%$ with shallow GNNs to $88\%$ with hundred- and thousand-layer GNNs, by capturing distant features. However, enforcing DP in multi-layer GNNs is particularly challenging, as these GNN models aggregate node embeddings over deeper layers and broader neighborhoods. Current research still lacks an effective solution to injecting DP noise to multi-layer GNNs with privacy-utility balance.
@@
Differentially private GNNs. Graphs consist of edges and nodes. Corresponding to instance-level DP [csf/mironov2017renyi, ccs/AbadiCGMMT016, nips/0001S22], the “instance” of graphs can be an edge or a node, naturally called edge-level DP (EDP) and node-level DP (NDP). GNNs have emerged as a key approach for applications over graph-structured data, such as intrusion detection [eurosp/ApruzzeseLS23, sp/SommerP10], social recommendation [icde/XiaSH0XP23], and drug discovery [kdd/ZhongM23]. Sharing trained GNN model can lead to privacy risks [uss/000100S022, ccs/0001MMBS22, corr/abs-2102-05429], typically membership inference attack (MIA) [sp/CarliniCN0TT22, ccs/0001MMBS22]. MIA stems from “overfitting”, where models can memorize training memberships [zhang2021understanding], either an edge or a node. Consequently, GNNs can leak sensitive information about their edge- or node-level neighbors.
@@
To address these risks, existing research works [arxiv/sajadmanesh2020differential, arxiv/Wuprivacy23, nips/chien2023differentially, arxiv/DaigavaneNode21, corr/abs-2109-08907, uss/sajadmanesh2023gap] have integrating DP with GNNs to achieve EDP and NDP. One research direction is to utilize graph perturbation (e.g., LPGNet [ccs/KolluriBHS22] and LapGraph [sp/0011L0022]) through a randomized response mechanism and adding discrete DP noise to the adjacency matrix. Then, the perturbed graph is passed to GNNs for subsequent training tasks, where the graph perturbation is required only once and also irrelevant to the GNN architectures. However, the GNN model utility is low when being trained over a perturbed graph when the privacy budget is tight, for example, $40\%$ accuracy of $\epsilon=1,2,3,4$ reported in LPGNet [ccs/KolluriBHS22].
@@
To improve utility, perturbed message-passing mechanism (PMP) [uss/sajadmanesh2023gap] has been proposed by adding the calibrated Gaussian noise to the message-passing layer, and DPDGC perturbs the decoupled graph convolution [nips/chien2023differentially]. As PMP realizes a better trade-off of privacy and utility, our work extends the research line of PMP. Table I presents a comprehensive comparison. Albeit their efforts on EDP and NDP, leveraging the contractive hidden node embeddings in private GNNs for amplifying privacy remains an underexplored avenue; thus, $\mathsf{CARIBOU}$ fills this gap. More related works are detailed in Appendix A-I.

### v3

@@ III-A2 Empirical (Counter-Intuitive) Observation: Deeper GNNs May Enhance Privacy
Recent work by [81] revealed a counter-intuitive phenomenon: deeper GNNs empirically exhibit lower vulnerability to membership and link inference attacks. This challenges standard privacy composition analysis, which suggests privacy risk increases with model depth due to multiple queries of the graph data. The insight stems from the over-smoothing [^1] Incidentally, another similar phrase is “over-fitting”, which refers to a model performing well on training data but poorly on unseen data due to memorization. We emphasize that they are different concepts to avoid confusion. [8] phenomenon, where node representations become indistinguishably homogeneous phenomenon, where node representations converge toward similar values as depth increases, making it inherently difficult for adversaries to distinguish individual nodes or infer sensitive relationships.
@@ III-B Core Idea for Convergent Privacy
Designing Private GNNs with Convergent Privacy. We aim to leverage the privacy amplification properties of contractive map to design a new framework for private GNNs. This framework is inspired by the recent advances in differentially private gradient descent (DP-GD) [14, 6], which has shown that the privacy cost can converge to a finite value even with arbitrarily many iterations. Motivated by it, we aim to translate the advanced privacy analysis techniques from DP-GD to GNNs. We observe that perturbed message passing (Equation 2) in GNNs follows a strikingly mathematical parallel pattern as DP-GD: This parallel structure enables to leverage the insight of convergent privacy analysis in the context of GNNs, provided we ensure two critical conditions:
@@ IV Private GNNs with Contractiveness
Security Model. Aggregation-based GNNs such as GCN [35], GCN-II [11], and SAGE [27] reveal only the final node embeddings, keeping intermediate embeddings hidden. The final node embeddings learned in this DP manner can be subsequently applied to downstream tasks such as node classification [75], link prediction [56], and graph classification [40]. Since DP introduces utility loss, GAP [59] seeks to improve model utility by concatenating all intermediate node embeddings from the $K$ layers. However, it causes the noise scale $\sigma$ to grow linearly as $O(\sqrt{{K}/{\epsilon^{2}}})$, and this privacy bound is loose. This work takes a step further beyond GAP, removing the assumption that all intermediate node embeddings must be revealed under a more realistic security model [91].
@@ IV-A Contractive Aggregation Module
Contractive operations in GNNs, such as the graph convolution (GConv) layer [36], inherently reduce privacy risks by mitigating the memorization of GNN parameters through the over-smoothing phenomenon [8]. This property aligns well with the need for $K$-layer aggregation in long-range graphs, where a target node aggregates embeddings from distant source nodes. However, directly stacking $K$ GConv layers, as shown in Figure 3, is suboptimal due to limited expressive power [11] and heightened sensitivity to DP noise [79].
@@
Comparison with existing design. GAP [59] enhances expressiveness by connecting all intermediate embeddings but lacks contractiveness, compromising utility and privacy. In contrast, CGL introduces adjustable coefficients, achieving a balanced trade-off between privacy, utility, and generalizability. Table I summarizes the general design characteristics between GAP and $\mathsf{CARIBOU}$.
@@ IV-C Privacy Auditing Module
Message passing mechanisms integrate graph structures by recursively aggregating and transforming node features from neighbors. Membership Inference Attacks (MIA) on graph examine the vulnerability of message passing to infer whether specific nodes or edges were part of a GNN’s training set [53]. The adversaries exploit black-box access to the GNN, querying it with selected data and analyzing outputs (e.g., class probabilities) to infer membership. To align theoretical privacy guarantees with practical deployments, we propose an empirical auditing module. This module simulates an adversary to evaluate GNN privacy before deployment. Extending Carlini et al. [7]’s MIA framework to graphs, we define a graph-specific privacy auditing game (Definition 7) and implement two real-world attacks [78, 53] for privacy verification.
@@ IV-D Putting it Together
To improve classification accuracy, $\mathsf{CARIBOU}$ adopts a head MLP architecture proposed by Sajadmanesh et al. [59] as the CM. This design ensures that the CM effectively combines the information from both feature sets, enabling robust node classification. Furthermore, the CM guarantees efficient training by leveraging the graph-agnostic features ${\bm{X}}^{(0)}$, ensuring a lower-bound performance even in scenarios where graph topology is unavailable.
@@ V Convergent Privacy Analysis
In this section, we present a convergent privacy analysis for perturbed message-passing GNNs with respect to the number of hops. In Appendix A-A, we review the standard privacy analysis for a one-layer perturbed message-passing GNN, and then illustrate that the privacy cost grows linearly with the number of layers under standard composition theorems. We then shift our focus to a convergent privacy analysis for perturbed GNNs whose message-passing layers are contractive. In particular, we draw upon the framework of contractive noisy iteration (CNI) from [6], recasting the multi-layer perturbed GNN as a CNI process. Our analysis reveals that, under hidden intermediate states and contractive message-passing layers, the privacy cost converges as the number of hops increases. Finally, we specialize this result to our proposed $\mathsf{CARIBOU}$. We show that $\mathsf{CARIBOU}$’s message-passing operation is contractive, derive its sensitivity, and thereby establish concrete bounds on both edge-level and node-level differential privacy for arbitrarily many hops. All proofs are deferred to the Appendices A-B,A-C.
@@ V-A Contractive Noisy Iteration and Convergent Privacy
Many GNNs, such as GCN, use "mean-type" aggregation, mixing a node’s representation with that of its neighbors. Intuitively, iterative mixing could "amplify" privacy, but existing analyses yield only linear compositions. Our key observation is that perturbed message passing in a contractive GNN layer behaves analogously to noisy gradient descent or noisy iterative maps [14, 4, 3], where recent work has demonstrated privacy amplification via iteration.
Below, we introduce the framework of contractive noisy iteration (CNI) from [6], and the meta theorem proved by [6] that provides a tight privacy guarantee for CNI processes.
Definition 8 (Contractive Noisy Iteration (CNI)[6, Definition 3.1]).
@@
A special case of CNI considered in [6] is the noisy gradient descent, where the contractive functions are the gradient update steps for some fixed loss function $f$, and the noise distribution $\xi_{k}$ is Gaussian. This situation is similar to the perturbed message passing mechanism in a contractive GNN layer, where the contractive function $\phi_{k}\equiv\mathrm{MP}_{G}$ is the message passing operation for a fixed graph $\bm{G}$}.
The complete privacy analysis of CNI processes in [6] involves concepts of trade-off functions (Definition 9) and Gaussian differential privacy (GDP, Definition 10), which we briefly review in Appendix A. GDP can be converted to the more familiar Rényi Differential Privacy (RDP) framework:
Lemma 1 (GDP Implies RDP [6, Lemma A.4]).
@@
With this connection established, we now state the key meta theorem from [6] that analyzes the CNI process and provides a tight privacy guarantee.
Theorem 2 (Meta Theorem on CNI [6, Theorem C.5]).
@@
The theorem above slightly generalizes the original result from [6] by relaxing the Lipschitz condition to require $\gamma<1$ only for $k\geq 2$ rather than for all iterations. This relaxation is critical for analyzing our $\mathsf{CARIBOU}$ architecture, where the first message-passing layer includes the residual term $\beta{\bm{X}}^{(0)}$, potentially making it non-contractive while subsequent layers remain contractive. The proof extends the original argument by carefully tracking the influence of the first layer on the privacy guarantee. For completeness, we provide the full proof in Appendices A-B,A-C.
@@
Theorem 3 ([6, Theorem 4.2] adapted for GNNs).
@@ VI Experimental Evaluation
Datasets. $\mathsf{CARIBOU}$ was tested over nine graph datasets. Five of the datasets have been broadly used to evaluate GNN methods, including Photo and Computers [62], Cora and PubMed [61], Facebook [71]. We also adjusted the synthetic chain-structured dataset developed under [25] into various scales, termed Chain-S, Chain-M, Chain-L and Chain-X. The chain-structured dataset has been used to understand the relations between privacy/utility and hops, as described in Section III-A. It is considered as an important benchmark to assist the development of long-range interaction GNNs by the ML community [18]. More details on datasets, model configurations, and privacy configurations are specialized in Appendix A-D1.
Baselines. We compare $\mathsf{CARIBOU}$ with multiple baselines [13, 59, 77, 15, 38] about edge-level private GNNs and vanilla GNNs. To our best knowledge, GAP [59] and DPDGC [13] are the strongest baselines for perturbed message passing under Gaussian mechanisms in both edge/node-level DP GNNs. In addition, we consider another research direction, i.e., first perturbing the graph through randomized response and then training GNNs over the perturbed graph PertGraph [77]. For a comprehensive evaluation, we adopt both research lines of works as our baselines. MLP is a baseline commonly compared with GNNs to demonstrate how well GNNs utilize graph structures, as it provides a reference counterpart to which GNNs learn the only node features without graph topology. More details of the baseline methods are provided in Appendix A-D3.
@@ VI-B Privacy Auditing
Following the PAM outlined in Section IV-C, we focus on black-box, membership-based privacy audits that match the theoretical guarantees of our DP mechanisms and those of prior perturbed message-passing methods [13, 59]. Under this threat model, LinkTeller [78] and G-MIA [53] are canonical and complementary: LinkTeller targets edge-level membership by asking whether a specific edge exists in the training graph; G-MIA targets node-level membership by deciding whether a node and its connected edges were used during training. We adopt G-MIA’s attacking settings of TSTF, where models have been trained on subgraph and tested on full graph. The adversary knows the whole graph $\bm{G}$} and all edges contained in $\bm{G}$} but has no access to the subgraphs used for early training. Both attacks (i) are specifically designed for GNNs, (ii) operate in the transductive setting considered in our analysis, (iii) require only query access to GNN models, and (iv) are publicly available and already used to evaluate DP-GNN defenses. This makes them ideal choices as mechanism-level auditing tools in $\mathsf{CARIBOU}$.
In Table XI (in Appendix), we report the AUC (Area Under the Curve) score about the attack effectiveness, when $\mathsf{CARIBOU}$ is being attacked. AUC is a major metric to evaluate the membership inference attack [78, 28]. Specifically, AUC measures true positive rate against the false positive rate on various classification thresholds, and a score of $0.5$ suggesting random guessing. We found $\mathsf{CARIBOU}$ is very effective against LinkTeller, by dropping the attack AUC from between 0.86 to 0.998 across all standard datasets ($\epsilon=\inf$) to less than 0.500 ($\epsilon$ ranges from 1 to 32). The similar effect was also observed by Wu et al. (e.g., less than 0.5 attack precision for 3-layer GCN for high density belief, shown in Table IX) [78] and Tang et al. (e.g., less than 0.5 attack AUC sometimes in Figure 10) [67]. For G-MIA, its AUC on the 5 datasets are already lower than LinkTeller by a notable margin when $\epsilon=\inf$ (between 0.567 to 0.702 for the 5 datasets), so the impact of $\mathsf{CARIBOU}$ is relatively small. But we observe on Cora, $\mathsf{CARIBOU}$ is able to drop AUC from 0.645 to 0.500.
@@ VIII Related Works
Multi-layer GNNs. Recent literature shows that multi-layer GNNs hold significant potential for modeling long-range dependencies and complex relational structures crucial for many real-world applications. Node labels and attributes may depend on distant nodes, necessitating the aggregation of information over larger receptive fields [85] through multi-layer GNNs. Notably, Li et al. [39] demonstrated, through the 1000-layer GNN, that increasing the network depth attains substantial gains in accuracy, e.g., from $72\%$ with shallow GNNs to $88\%$ with hundred- and thousand-layer GNNs, by capturing distant features. However, enforcing DP in multi-layer GNNs is particularly challenging, as these GNN models aggregate node embeddings over deeper layers and broader neighborhoods. Current research still lacks an effective solution to injecting DP noise to multi-layer GNNs with privacy-utility balance.
@@
Differentially private GNNs. Graphs consist of edges and nodes. Corresponding to instance-level DP [49, 1, 91], the “instance” of graphs can be an edge or a node, naturally called edge-level DP (EDP) and node-level DP (NDP). GNNs have emerged as a key approach for applications over graph-structured data, such as intrusion detection [5, 63], social recommendation [86], and drug discovery [101]. Sharing trained GNN model can lead to privacy risks [99, 90, 29], typically membership inference attack (MIA) [7, 90]. MIA stems from “overfitting”, where models can memorize training memberships [94], either an edge or a node. Consequently, GNNs can leak sensitive information about their edge- or node-level neighbors.
@@
To address these risks, existing research works [57, 82, 13, 15, 52, 59] have integrating DP with GNNs to achieve EDP and NDP. One research direction is to utilize graph perturbation (e.g., LPGNet [38] and LapGraph [77]) through a randomized response mechanism and adding discrete DP noise to the adjacency matrix. Then, the perturbed graph is passed to GNNs for subsequent training tasks, where the graph perturbation is required only once and also irrelevant to the GNN architectures. However, the GNN model utility is low when being trained over a perturbed graph when the privacy budget is tight, for example, $40\%$ accuracy of $\epsilon=1,2,3,4$ reported in LPGNet [38].
@@
To improve utility, perturbed message-passing mechanism (PMP) [59] has been proposed by adding the calibrated Gaussian noise to the message-passing layer, and DPDGC perturbs the decoupled graph convolution [13]. As PMP realizes a better trade-off of privacy and utility, our work extends the research line of PMP. Table I presents a comprehensive comparison. Albeit their efforts on EDP and NDP, leveraging the contractive hidden node embeddings in private GNNs for amplifying privacy remains an underexplored avenue; thus, $\mathsf{CARIBOU}$ fills this gap. More related works are detailed in Appendix A-I.